# ReCAP: Recursive Cross Attention Network for Pseudo-Label Generation in Robotic Surgical Skill Assessment

### v2

@@ Abstract
In surgical skill assessment, Objective Structured Assessments of Technical Skills (OSATS scores) and the Global Rating Scale (GRS) are established tools for evaluating the performance of surgeons during training. These metrics, coupled with feedback on their performance, enable surgeons to improve and achieve standards of practice. Recent studies on the open-source dataset JIGSAW, which contains both GRS and OSATS labels, have focused on regressing GRS scores from kinematic signals, video data, or a combination of both. In this paper, we argue that regressing the GRS score, a unitless value, by itself is too restrictive, and variations throughout the surgical trial do not hold significant clinical meaning. To address this gap, we developed a recurrent transformer model that outputs the surgeon’s performance throughout their training session by relating the model’s hidden states to five OSATS scores derived from kinematic signals. These scores are averaged and aggregated to produce a GRS prediction, enabling assessment of the model’s performance against the state-of-the-art (SOTA). We report Spearman’s Correlation Coefficient (SCC), demonstrating that our model outperforms SOTA models for all tasks, except for Suturing under the leave-one-subject-out (LOSO) scheme (SCC 0.68-0.89), while achieving comparable performance for suturing and across tasks under the leave-one-user-out (LOUO) scheme (SCC 0.45-0.68) and beating SOTA for Needle Passing (0.69). We argue that relating final OSATS scores to short instances throughout a surgeon’s procedure is more clinically meaningful than a single GRS score. This approach also allows us to translate quantitative predictions into qualitative feedback, which is crucial for any automated surgical skill assessment pipeline. A senior surgeon validated our model’s behaviour and agreed with the semi-supervised predictions 77 % ($p=0.006$) of the time.
@@ I Introduction
Recognizing the need for a standardised framework for surgical skill assessment, the Objective Structured Assessment of Technical Skills (OSATS) [19] score was developed. This Likert-style scale assesses various core components in surgery. The summation of individual OSATS scores yields a Global Rating Score (GRS), utilised to assess surgical trainees’ performance across multiple procedures. OSATS provides quantitative and qualitative performance analysis, irrespective of the level of expertise.
While these metrics are able to capture an overall picture of performance, they rely on the assessors’ feedback, typically senior clinicians [2], to capture intraoperative performance [3] during surgical trials.
@@
The JHU-ISI Gesture and Skill Assessment Working Set (JIGSAW) dataset [30] is currently used as the benchmark dataset for surgical skill assessment. While recent studies on the JIGSAW dataset have shifted towards video-centric approaches, the field of surgical skill assessment still relies on kinematic data [23, 25, 26, 24]. Kinematic data offer standardisation across datasets whereas video data is prone to considerable variations. Moreover, kinematic data imposes lower computational demands, allowing for rapid and efficient development processes. Lastly, we argue that the use of kinematic data instead of video is more ethical. Considering these factors, we direct our focus towards kinematic data.
@@
While existing ML models excel at categorising surgical expertise [14, 8] and predicting overall performance metrics like the GRS  [17, 18, 20, 8, 14] on the JIGSAW dataset  [30], they often fall short in capturing nuanced variations throughout a surgical procedure. Recent research efforts have aimed to address this gap by exploring models that can track changes in GRS or OSATS scores over the course of a procedure [14, 22, 12, 20, 8, 13, 17]. Yet, many of these approaches either increase clinician workload [12, 20] by requiring additional labels or lack meaningful validation [8, 13, 17]. It’s also important to note that while the interpretability mechanisms proposed in these models are technically sophisticated, e.g. counterfactual explanations [31], they often don’t translate well into clinical feedback [16, 15]. Wang et al. [20] proposed a framework that identifies problematic events throughout a surgical trial. They employed a supervised recurrent network to aggregate intermediate Global Rating Scale (GRS) scores into a final score. However, the reliance on more granular and additional labels along with the inherent lack of descriptive qualities in the GRS score are notable drawbacks of their method. Anastasiou et al. [17] adopted a contrastive approach for regressing the GRS score, facilitating a straightforward interpretation of performance relative to a reference. However, deviations from this reference lack translation into actionable feedback or clinical interpretability. A deviation of 10 points from a reference doesn’t allow the trainee to know exactly where he underperformed when the score he receives is an aggregate of 6 distinct surgical skills. Zia et al. [8] investigated how specific segments of the input data, through hand-crafted features, influence OSATS score predictions. While this approach yields more meaningful insights into the variations of performance, they do not offer any actionable information to the clinician. The change in predictions cannot be mapped in a one-to-one fashion to the qualitative description the OSATS give. Similarly, Fawaz et al. [14] extracted performance variations by regressing OSATS scores, whereby GRAD-CAM [27] is used to pinpoint positive and negative contributions to the final output. Nevertheless, both methods fall short of providing clinically comprehensive feedback.
@@
    The building blocks for task agnostic models to relate segment-level OSATS predictions to qualitative feedback.
@@ II-A Problem Formulation
An input signal $X_{i}\in\mathbb{R}^{D\times T_{i}}$, of feature size D and length $T_{i}$, is divided into equal segments $x^{s}_{i}$ of size $L$ ($x^{s}_{i}\in\mathbb{R}^{D\times L}$): $\{x^{1}_{i},x^{2}_{i},\dots,x^{s}_{i}\}\in X_{i}$ where $S_{i}$ is the total number of segments, i.e. $S_{i}=\frac{T_{i}}{L}$ for given signal index i. For simplicity, in the rest of this paper, we omit $i$ and use s as a subscript to refer to different segments within a trial $i$, i.e. $x^{s}_{i}\to x_{s}$. We fit a function $F$ to map $X$ to the label space $\mathcal{Y}:(F:X\to\vec{y})$:
@@ II-B Model Overview
Our model processes segments of a kinematic signal recurrently by taking two inputs: the previous hidden state of the recurrent network, $z_{s-1}\in\mathbb{R}^{D\times L}$, and the current segment-level kinematic signal, $x_{s}$ (Fig 1a). The two inputs are fused into the current hidden state, $z_{s}=h(x_{s},z_{s-1})$, through the model backbone $h$ (Fig. 1b). We initialise $z_{0}$ as a zero-filled tensor. Each hidden state is then passed to six classification heads, $c_{n}$, giving $f_{n}(x_{s})=c_{n}(h(x_{s},z_{s-1}))$. The output of our model is a final OSATS, the average of all segment-level OSATS:
@@
Figure 1: ReCAP Architecture Overview: (a) A kinematic signal is split into segments $x_{s}$ of size $L$ and used as inputs to our backbone $h$ recurrently. The previous state $z_{s-1}$ is also passed as an input and fused to $x_{s}$ to produce $z_{s}$ (c) This fusion is performed by three fusion modules consisting of self- and cross-attention blocks. (b) The current hidden state $z_{s}$ is given as an input to five classification heads $c_{n}$ to predict the respective OSATS scores . $y_{1s}$ to $y_{5s}$ corresponds to the respective OSATS category prediction of Time and Motion, Flow of Operation, Suture/Needle Handling, Respect of Tissue, and Overall Performance at segment position $s$
The classification heads $c_{n}$ are five MLPs classifying the hidden state $z_{s}$ into segment-level OSATS predictions $\hat{y}_{ns}=c_{n}(z_{s})$. Each MLP layer consists of batch normalisation, ReLU activation function, and fully connected layers.
Loss: ReCAP is trained end-to-end using cross-entropy losses. Each cross-entropy loss is applied to the average of the classification head segment predictions $\hat{y_{n}}=\frac{1}{S}\sum\hat{y}_{ns}$ for a given OSATS category label $y_{n}$. An $L2$ penalty term is added to regularize the network and help with generalisation. Our final loss is expressed as:
@@
TABLE I: Performance comparison of GRS score on JIGSAW for the LOSO cross-validation scheme of models trained on independent and across tasks. K: Kinematic, V Video. *: Results from training on the three domains.
Task & Scheme
@@
II-C Experimental Design
Dataset: We evaluated our model on the JIGSAWS dataset [30]. This dataset consists of video and kinematics data generated by eight clinicians evaluated on three distinct tasks, namely needle passing (NP), suturing (SU), and knot-tying (KT). Altogether, there are 39, 28, and 36 labelled data samples for SU, NP, and KT, respectively. The labels are comprised of six OSATS(1-5) and one GRS(6-35) score the aggregate of all OSATS. It is worth noting that the score 5 only appears in the suturing task. The OSATS include 1) respect for tissue, 2) suture/needle handling, 3) time and motion, 4) flow of operation, 5) overall performance, and 6) quality of the final product. In this study, we only use kinematic data.
Cross-validation: Following the JIGSAW cross-validation framework, we evaluate our method using Leave-One-Supertrial-Out (LOSO) whereby the i-th trial performed by the surgeons are left out as the validation set. Another cross-validation scheme Leave-One-User-Out (LOUO) is not considered in this work. The lack of literature reporting OSATS performance on this scheme and very imbalanced training/validation folds were reasons enough. Recent literature [32, 17] on the Jigsaw dataset seems to indicate that no testing fold is used to assess performance.
Performance Metric and Reporting: Similar to relevant work [8, 6, 7, 28], we evaluate our method using Spearman’s Correlation Coefficient (SCC) $\rho$ to compare the predicted ranked GRS score with the ground truth (Eq. 5). We report SCC averaged across folds for the LOSO cross-validation scheme under the same training parameters and at the same training epoch. The intermediate OSATS are averaged into a signal-level OSATS after processing the whole kinematic sample. The final six OSATS are then summed to give a video-level GRS score. Note that the predicted GRS score, $\hat{Y}=\sum_{n=0}^{6}\hat{y_{n}}$, is only used to assess model performance, but is not directly learned from our model:
The mean average error (MAE) is also reported at the best-performing epoch for this specific metric. To our knowledge, Benmansour et al [32] are the only recent work reporting OSATS-specific performance under the LOSO validation scheme. OSATS performance is reported at the same epoch used to report GRS performance under the same training parameters. OSATS SCC was averaged across 10 epochs to mitigate the intrinsic variability coming from the short range of labels. OSATS are learned and regressed by our model.
Data Augmentation: Two augmentation techniques were added to the kinematic signals to improve generalisation: 1) Gaussian noise based on the standard deviation of the signal, and 2) flipping, i.e. reversing the signal. Augmentations were done at a rate of $50\%$. Label smoothing was also performed at 30%.
Validation of Model Behaviour: To validate our model’s ability to generate interim OSATS scores, we asked a consultant surgeon in endoscopic interventions to agree or disagree with the model’s intermediate predictions. Every 75 frames were assigned a generated pseudo-label. The label is shown on the screen during viewing. Similar to Wang et al.’s framework [20], the surgeon was aware of the ground truth assigned to the video. The predicted OSATS overall performance score was divided into three categories: poor (1-2), average (3), and good (4-5). We randomly introduce noise to the model’s predictions to mitigate potential bias without informing the surgeon. We then present these predictions and capture agreement or disagreement at the segment level when playing the video sequentially.
TABLE II: Perofrmance for OSATS scores, where the $\rho Osats$ is the average across the 6 scores under LOSO scheme. *: results from training across the 3 tasks.
KT NP SU AT
Apen[8] 0.66 0.45 0.59 0.57
[…]
III Results
TABLE III: Performance for OSATS scores, where the $\rho Osats$ are reported for RT: Respect for tissue. TM: Time and Motion. OP: Overall Performance under the LOSO in the KT task.
CNN + Bilstm     [32] ReCAP
[…]
RT 0.83 0.49 0.46 0.92/0.78 0.75/0.43 0.78/0.52 0.56
TM 0.87 0.85 0.68 0.95/0.8 0.91/0.72 0.84/0.60 0.62
OP 0.89 0.58 0.71 0.9/0.79 0.42/0.23 0.69/0.5 0.65
SNH 0.82 0.79 0.75 0.84/0.61 0.91/0.69 0.88/0.78 0.65
FO 0.76 0.58 0.62 0.78/0.63 0.66/0.45 0.89/0.66 0.64
QFP 0.75 0.31 0.67 0.85/0.59 0.56/0.22 0.91/0.64 0.62
Mean 0.82 0.60 0.65 0.87/0.70 0.70/0.46 0.83/0.62 0.62
We report the performance of our model against previous work that uses kinematic data or video data and report GRS performance (Table II-B). Although we don’t regress the GRS’s most recent work only report on it. To allow for easier comparison we use the GRS as a performance proxy. The model outperforms all methods using kinematic data and achieves competitive performance against models using video (Table II-B). When looking at the performance of our model in predicting OSATS scores under the LOSO validation scheme, we underperform only in NP (Table II). The CNN+Bilstm [32] only reports the best-performing fold. We also report the average across the 5 folds. We only underperform for the OSATS of Overall Performance in needle passing.
We performed two ablation studies to understand the contribution of each element of the model on the performance. The introduced guassian noise and flipping had very little effect on the performance of the model. However the the flipping does allow for the model to be time invariant. We see that the pseudo-label drastically improves performance, especially for the two tasks, NP and SU, with the most class imbalance [29].
@@ III Results
To validate the weakly supervised outputs, 9 videos were reviewed by a consultant surgeon, where each 75 frames (the segment length) had an assigned OSATS pseudo-label. The selected videos regrouped three levels of expertise (novice, intermediate, expert) across the three tasks. Two of those videos were shown with randomly generated predictions. We found that the clinician agreed 69% of the time when shown random noise while agreeing 77% of the time when shown our model’s predictions. A one-tailed binomial test between the two distributions indicates a statistically significant difference between the agreements (p=0.006).

### v3

@@ Abstract
In surgical skill assessment, the Objective Structured Assessments of Technical Skills (OSATS) and Global Rating Scale (GRS) are well-established tools for evaluating surgeons during training. These metrics, along with performance feedback, help surgeons improve and reach practice standards. Recent research on the open-source JIGSAWS dataset, which includes both GRS and OSATS labels, has focused on regressing GRS scores from kinematic data, video, or their combination. However, we argue that regressing GRS alone is limiting, as it aggregates OSATS scores and overlooks clinically meaningful variations during a surgical trial.
To address this, we developed a recurrent transformer model that tracks a surgeon’s performance throughout a session by mapping hidden states to six OSATS, derived from kinematic data, using a clinically motivated objective function. These OSATS scores are averaged to predict GRS, allowing us to compare our model’s performance against state-of-the-art (SOTA) methods. We report Spearman’s Correlation Coefficients (SCC) demonstrating that our model outperforms SOTA using kinematic data (SCC 0.83-0.88), and matches performance with video-based models.
Our model also surpasses SOTA in most tasks for average OSATS predictions (SCC 0.46-0.70) and specific OSATS (SCC 0.56-0.95). The generation of pseudo-labels at the segment level translates quantitative predictions into qualitative feedback, vital for automated surgical skill assessment pipelines. A senior surgeon validated our model’s outputs, agreeing with 77 % of the weakly-supervised predictions $p=0.006$.
@@ I Introduction
Recognizing the need for a standardised framework for surgical skill assessment, the Objective Structured Assessment of Technical Skills (OSATS) [19] score was developed. This Likert-style scale assesses various core components in surgery. The summation of individual OSATS scores yields a Global Rating Score (GRS), utilised to assess surgical trainees’ performance across multiple procedures. OSATS provides quantitative and qualitative performance analysis, irrespective of the level of expertise. While these metrics are able to capture an overall picture of performance, they rely on the assessors’ feedback, typically senior clinicians [2], to capture intraoperative performance [3] during surgical trials.
@@
The JHU-ISI Gesture and Skill Assessment Working Set (JIGSAWS ) dataset [30] is currently used as the benchmark dataset for surgical skill assessment. While recent studies on the JIGSAWS dataset have shifted towards video-centric approaches, the field of surgical skill assessment still relies on kinematic data [23, 25, 26, 24]. Kinematic data offers standardisation across datasets whereas video data is prone to considerable variations. Moreover, kinematic data imposes lower computational demands, allowing for rapid and efficient development processes. Lastly, we argue that the use of kinematic data is more ethical. Considering these factors, we direct our focus towards kinematic data.
@@
While existing ML models excel at categorising surgical expertise [14, 8] and predicting overall performance metrics like the GRS  [17, 18, 20, 8, 14] on the JIGSAWS dataset  [30], they often fall short in capturing nuanced variations throughout a surgical procedure. Recent research efforts have aimed to address this gap by exploring models that can track changes in GRS or OSATS scores over the course of a procedure [14, 22, 12, 20, 8, 13, 17]. Yet, many of these approaches either increase clinician workload [12, 20] by requiring additional labels or lack meaningful validation [8, 13, 17]. It’s also important to note that while the interpretability mechanisms proposed in these models are technically sophisticated, e.g. counterfactual explanations [31], they often don’t translate well into clinical feedback [16, 15]. Wang et al. [20] proposed a framework that identifies problematic events throughout a surgical trial. They employed a supervised recurrent network to aggregate intermediate Global Rating Scale (GRS) scores into a final score. However, the reliance on more granular and additional labels along with the inherent lack of descriptive qualities in the GRS score are notable drawbacks of their method. Anastasiou et al. [17] adopted a contrastive approach for regressing the GRS score, facilitating a straightforward interpretation of performance relative to a reference. However, deviations from this reference lack translation into actionable feedback or clinical interpretability. A deviation of 10 points from a reference doesn’t allow the trainee to know exactly where they underperformed as the score received is an aggregate of six distinct surgical skills. Zia et al. [8] investigated how specific segments of the input data, through hand-crafted features, influence OSATS score predictions. While this approach yields more meaningful insights into the variations of performance, they do not offer any actionable information to the clinician. The change in predictions cannot be mapped in a one-to-one fashion to the qualitative description the OSATS give. Similarly, Fawaz et al. [14] extracted performance variations by regressing OSATS scores, whereby GRAD-CAM [27] is used to pinpoint positive and negative contributions to the final output. Nevertheless, both methods fall short of providing clinically comprehensive feedback.
@@
    The building blocks for task agnostic models to relate segment-level OSATS predictions to qualitative feedback through a clincially motivated objective function.
@@ II-A Problem Formulation
An input signal $X_{i}\in\mathbb{R}^{D\times T_{i}}$, of feature size D and length $T_{i}$, is divided into equal segments $x^{s}_{i}$ of size $L$ ($x^{s}_{i}\in\mathbb{R}^{D\times L}$): $\{x^{1}_{i},x^{2}_{i},\dots,x^{s}_{i}\}\in X_{i}$ where $S_{i}$ is the total number of segments, i.e. $S_{i}=\frac{T_{i}}{L}$ for a given signal i. For simplicity, in the rest of this paper, we omit $i$ and use s as a subscript to refer to different segments within a trial $i$, i.e. $x^{s}_{i}\to x_{s}$. We fit a function $F$ to map $X$ to the label space $\mathcal{Y}:(F:X\to\vec{y})$:
@@ II-B Model Overview
Our model recurrently processes segments of a kinematic signal by taking two inputs: the previous hidden state of the recurrent network, $z_{s-1}\in\mathbb{R}^{D\times L}$, and the current segment-level kinematic signal, $x_{s}$ (Fig 1a). The two inputs are fused into the current hidden state, $z_{s}=h(x_{s},z_{s-1})$, through the model backbone $h$ (Fig. 1b). We initialise $z_{0}$ as a zero-filled tensor. Each hidden state is then passed to six classification heads, $c_{n}$, giving $f_{n}(x_{s})=c_{n}(h(x_{s},z_{s-1}))$. The output of our model is a final OSATS, the average of all segment-level OSATS:
@@
Figure 1: ReCAP Architecture Overview: A kinematic signal is split into segments $x_{s}$ of size $L$ and used as inputs to our backbone $h$ recurrently. The previous state $z_{s-1}$ is also passed as an input and fused to $x_{s}$ to produce $z_{s}$. This fusion is performed by a fusion module consisting of self- and cross-attention blocks. The current hidden state $z_{s}$ is given as an input to six classification heads $c_{n}$ to predict the respective OSATS. $\hat{y}_{1s}$ to $\hat{y}_{6s}$ corresponds to the respective OSATS category prediction at segment position $s$
The classification heads $c_{n}$ are five multilayer perceptron (MLPs) classifying the hidden state $z_{s}$ into segment-level OSATS predictions $\hat{y}_{ns}=c_{n}(z_{s})$. Each MLP layer consists of batch normalisation, ReLU activation function, and fully connected layers.
Loss: ReCAP is trained end-to-end using cross-entropy losses. Each cross-entropy loss is applied to the average of the classification head segment predictions $\hat{y_{n}}=\frac{1}{S}\sum\hat{y}_{ns}$ for a given OSATS category label $y_{n}$. An $L2$ penalty term weighted by $\lambda$ is added to regularize the network and help with generalisation. Our final loss is expressed as:
@@
II-C Experimental Design
Dataset: We evaluated our model on the JIGSAWS dataset [30]. This dataset consists of video and kinematics data generated by eight clinicians evaluated on three distinct tasks, namely needle passing (NP), suturing (SU), and knot-tying (KT). Altogether, there are 39, 28, and 36 labelled data samples for SU, NP, and KT, respectively. The labels are comprised of six OSATS(1-5) and one GRS(6-35) score the aggregate of all OSATS. It is worth noticing that the dataset is biased towards lower OSATS as the score 5 only appears in the suturing task. The OSATS include 1) respect for tissue, 2) suture/needle handling, 3) time and motion, 4) flow of operation, 5) overall performance, and 6) quality of the final product. In this study, we only use kinematic data.
Cross-validation scheme: Following the JIGSAWS cross-validation framework, we evaluate our method using Leave-One-Supertrial-Out (LOSO) whereby the i-th trial performed by the surgeons are left out as the validation set. The cross-validation scheme Leave-One-User-Out (LOUO) is not considered in this work since there is a lack of literature reporting OSATS performance on this scheme due to the imbalanced training/validation folds. Literature [32, 17] on the JIGSAWS dataset seems to indicate that no testing fold is used to assess performance.
Evaluation: Similar to relevant work [8, 6, 7, 28], we evaluate our method using Spearman’s Correlation Coefficient (SCC) $\rho$ to compare the predicted ranked GRS score with the ground truth:
We report SCC averaged across folds for the LOSO cross-validation scheme under the same training parameters and at the same training epoch. The intermediate OSATS are averaged into a signal-level OSATS after processing the whole kinematic sample. The final six OSATS are then summed to give a video-level GRS score. Note that the predicted GRS score, $\hat{Y}=\sum_{n=0}^{6}\hat{y_{n}}$, is only used to assess model performance, but is not directly learned from our model.
The mean average error (MAE) is also reported at the best-performing epoch for this specific metric. To our knowledge, Benmansour et al [32] is the only recent work reporting OSATS-specific performance under the LOSO validation scheme. OSATS performance is reported at the same epoch used to report GRS performance under the same training parameters. OSATS SCC was averaged across 10 epochs. OSATS are learned and regressed by our model.
Data Augmentation: Two augmentation techniques were added to the kinematic signals to improve generalisation: 1) Gaussian noise based on the standard deviation of the signal, and 2) flipping, i.e. reversing the signal. Augmentations were done at a rate of $50\%$. Label smoothing and dropout was performed at 30%.
Validation of Model Behaviour: To validate our model’s ability to generate interim OSATS scores, we asked a consultant surgeon in endoscopic interventions to agree or disagree with the model’s intermediate predictions. Every 75 frames were assigned a generated pseudo-label. The label is shown on the screen during viewing. Similar to Wang et al.’s framework [20], the surgeon was aware of the ground truth assigned to the video (the assigned OSATSs). The predicted OSATS overall performance score was divided into three categories: poor (1-2), average (3), and good (4-5). We randomly generate some segment predictions in two videos to mitigate potential bias without informing the surgeon. We then present these predictions and capture agreement or disagreement at the segment level when playing the video sequentially.
TABLE I: Perofrmance for OSATS scores, where the $\rho Osats$ is the average across the 6 scores under LOSO scheme. *: results from training across the 3 tasks. Across Tasks (AT)
KT NP SU AT
Apen[8] 0.66 0.45 0.59 0.57
[…]
III Results
We report the performance of our model against previous work that uses kinematic data or video data and report GRS performance (Table III). Although we don’t regress the GRS’s most recent work only report on it. To allow for easier comparison we use the GRS as a performance proxy. The model outperforms all methods using kinematic data and achieves competitive performance against models using video (Table III).
TABLE II: OSATS performance,$\rho Osats$ is reported for RT: Respect for tissue, TM: Time and Motion, OP: Overall Performance, SNH: Suture and Needle Handling, FO: Flow of Operation, QFP: Quality of Final Product for the best and average fold performance. Across tasks(AT) is trained on the three tasks
CNN+Bilstm ReCAP
[…]
RT 0.83 0.49 0.46 0.92/0.78 0.75/0.43 0.78/0.52 0.56
TM 0.87 0.85 0.68 0.95/0.8 0.91/0.72 0.84/0.60 0.62
OP 0.89 0.58 0.71 0.9/0.79 0.42/0.23 0.69/0.5 0.65
SNH 0.82 0.79 0.75 0.84/0.61 0.91/0.69 0.88/0.78 0.65
FO 0.76 0.58 0.62 0.78/0.63 0.66/0.45 0.89/0.66 0.64
QFP 0.75 0.31 0.67 0.85/0.59 0.56/0.22 0.91/0.64 0.62
AVG 0.82 0.60 0.65 0.87/0.70 0.70/0.46 0.83/0.62 0.62
When looking at the performance of our model in predicting OSATS under the LOSO validation scheme, we underperform only in NP (Table I). The CNN+Bilstm [32] only reports the best-performing fold. We report the average across the 5 folds. We only underperform for the OSATS of Overall Performance in needle passing.
TABLE III: Performance comparison of GRS score on JIGSAWS trained on independent and across tasks. K: Kinematic, V: Video. *: Results from training on the three domains.
Task
@@ III Results
We performed two ablation studies to understand the contribution of the augmentation and the modified objective function. The introduced guassian noise and flipping had very little effect on the performance of the model. However the the flipping does allow for the model to be time invariant. We see that the pseudo-label drastically improves performance, especially for the two tasks, NP and SU, with the most class imbalance [29].
@@
To validate the weakly supervised outputs, 9 videos were reviewed by a consultant surgeon, where each 75 frames (the segment length) had an assigned OSATS pseudo-label. The selected videos regrouped three levels of expertise (novice, intermediate, expert) across the three tasks. Two of those videos were shown with the randomly generated predictions. We found that the clinician agreed 69% of the time when shown random noise while agreeing 77% of the time when shown our model’s predictions. A one-tailed binomial test between the two distributions indicates a statistically significant difference between the agreements (p=0.006).